# The power-spectrum tensor in steady-state systems and its role in quantum friction

### v1

@@ Abstract
For systems in equilibrium, quantum statistical physics provides a number of general theorems and relations that are not tied to specific microscopic models, one example being the fluctuation-dissipation theorem. Much less is known for non-equilibrium situations. In this work, we discuss certain properties of the power-spectrum tensor for systems in general steady-states, i.e. stationary states not necessarily corresponding to equilibrium configurations. In our analyses, we do not make any direct connection to specific models for the underlying microscopic dynamics and, therefore, our results can be applied to a large variety of systems. We also connect the power-spectrum tensor to other quantities that characterize these systems and, where appropriate, compare with the equilibrium counterparts. As an application, we consider the specific problem of quantum friction, where, at zero temperature, a contactless quantum-electrodynamic drag force acts on a particle that moves in close proximity to an arrangement of material bodies. Specifically, we show how the additional information about the system’s physics facilitates the derivation of more precise constraints on the power spectrum and its functional dependencies.
I Introduction
@@
Stochastic processes are omnipresent in engineering, physics and chemistry, so that a detailed understanding of their properties lies at the heart of many applications both, for fundamental and technological purposes. Specifically, the power spectral density is one of the most relevant quantities for studying noisy time-stationary systems: It describes how the strength (commonly the power) of different components of stochastic processes is distributed across different frequencies [1]. In stationary systems it is connected to the correlation function via the Wiener-Khinchin theorem [1], establishing that one quantity is the Fourier transform of the other. Whenever the stochastic process is multivariate or multidimensional, both the correlation and the power spectral density become tensorial in nature and we speak of correlation and power-spectrum tensors, respectively. In addition to the autocorrelations of each of the elements of the stochastic process, these quantities also contain the cross-correlations between them.
@@
In thermal equilibrium, a system’s power-spectrum tensor is connected to the system’s susceptibility tensor by way of the fluctuation-dissipation theorem (FDT) [2, 3]. The FDT is related to the detailed balance between the different energy-exchange processes that take place in equilibrium systems [4, 5]. Probably one of the most remarkable feature of this relation is its generality: As long as the system is in thermal equilibrium, the details of the underlying microscopic dynamics do not matter – whether the dynamics is linear or nonlinear, whether it is quantum or classical, whether it is Markovian or not [4, 1, 6, 7], the power spectrum is always proportional to the susceptibility via a known factor that depends on temperature. Therefore, it is not at all surprising that the FDT is applied in numerous and very diverse physical systems [8, 9, 10, 11, 12, 13, 14], sometimes even transcending its range of applicability. In fact, the FDT has been considered also for nonequilibrium systems for which, technically speaking, it becomes inaccurate [15, 16, 17, 18, 19, 20, 21, 22, 23]. Arguments for the validity of this approach have typically been based on the so-called local thermal equilibrium (LTE) approximation. Within the LTE approximation, it is assumed that the correlation length associated with the fluctuating dynamics is so short that the system’s spatially separated components are (locally) in thermal equilibrium with their immediate surroundings [24, 25, 26]. However, the quality of this approximation depends strongly on the specific system and its parameters, thus rendering it at least questionable in a number of cases [24, 26].
@@
The desire for deeper insights into nonequilibrium physics, also in connection to the hindsight provided by stochastic thermodynamics [27, 28], has recently motivated the research on more general fluctuation-dissipation and thermodynamic uncertainty relations [6, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], i.e. connections between different physical quantities and bounding inequalities with a broad range of validity, even beyond equilibrium. Within this context, nonequilbrium and the breaking of equilibrium relations has been investigated theoretically [39, 40, 41, 42, 43] and experimentally in systems such as polymers [44], colloidal gases [45], and spin glasses [46]. Particular attention has been devoted to systems in nonequilibrium steady-states (NESS) [47, 41, 5, 48, 49, 50, 51, 52, 30]. In such stationary states, a system’s stochastic properties are translation invariant with respect to time, despite the fact that the system is not in equilibrium. To exist, the NESS must thus be sustained by one or several external agents such as externally maintained temperature gradients, chemical potential differences, or even mechanical drives that act on the system. In general, this can break the aforementioned detailed balance and, therefore, lead to violations of the FDT [5].
@@
In this work, we also focus on a stationarity and discuss some properties of the power-spectrum tensor for a system that is characterized by being in a steady-state. While we address nonequilibrium dynamics, our approach also includes the equilibrium case as a special limit. Specifically, we do not make any detailed assumption about the microscopic dynamics of the system, thereby keeping the level of generality similar to that of the FDT. The stationarity of the NESS still allows to make certain general statements that can be related to the properties of the underlying stochastic dynamics.
@@
We further discuss the properties of the power-spectrum tensor in the case of the nonconservative viscous motion of a particle in the quantum electromagnetic vacuum [53, 54]. Specifically, we address quantum friction, i.e. the drag force acting on a neutral particle driven with constant velocity in the empty space close to a generic arrangement of bodies. Still keeping the microscopic dynamics unspecified, we show how additional information about the physics of the system strongly constrains the functional behavior and the structure of the power-spectrum tensor.
@@
The paper is organized as follows. In Sec. II, we consider the power-spectral density in the steady-state of a multidimensional stochastic process. In Sec. III, we specify our analysis to the case of quantum friction and show how the thus increased information about the system allows us to derive additional relations. To confirm our results and illustrate the advantages as well as the simplifications introduced by our general approach, we consider here an exactly solvable model for describing the internal dynamics of the particle. Finally, in Sec.IV we summarize and discuss our findings.
II The power-spectrum tensor in steady-state systems
@@
Let us consider a generic physical system described by the Hamiltonian $\hat{H}_{0}$. Our focus here is on a quantum system with a large number of degrees of freedom. It is thought, in general, as being composed of a collection of interacting subsystems: On its own each of the quantum subsystems is open, i.e. able to exchange energy and other physical quantities with the remaining parts of the system, representing its environment [55, 56]. Contrary to typical approaches, however, we will refrain from deriving any master equation, avoiding the Markov and other approximations which, depending on the specific circumstances, can be problematic for the description of quantum systems [4, 7, 57]. Indeed, except for the fact that the full system is in a unique stationary state described by the density matrix $\hat{\rho}$ with $[\hat{\rho},\hat{H}_{0}]=0$, we do not rely on any specific microscopic information about its dynamics. Again, we would like to note explicitly that $\hat{\rho}$ does not necessarily correspond to an equilibrium state.
@@
Let $\hat{d}_{i}$ be the components of a Hermitian vector-operator $\hat{\mathbf{d}}$. The time-evolution of each component, $\hat{d}_{i}(t)$, is induced according to $\hat{H}_{0}$. For our purposes, we consider a three-dimensional vector but generalizations to different dimensions are possible. The vector-operator $\hat{\mathbf{d}}$ either corresponds to a three-dimensional observable ($i=x,y,z$, for instance a particle’s electric dipole moment – see Sec. III), or a quantity that characterizes a multivariate stochastic process as for example in three-particle systems. For simplicity, in the subsequent discussion, we will refer to the former case but most of the results can be transferred to the latter case upon adequate replacements of the relevant quantities.
@@
Since the system is stationary, the expectation value of $\hat{\mathbf{d}}(t)$ over $\hat{\rho}$ is constant. Without loss of generality, we assume it to be zero [58]. i.e. $\langle\hat{d}_{i}(t)\rangle\equiv\mathrm{tr}[\hat{d}_{i}(t)\hat{\rho}]=\mathrm{tr}[\hat{d}_{i}\hat{\rho}]=0$, where ‘$\mathrm{tr}$’ indicates the trace over a generic set of basis vector-states of the system’s Hilbert space. Given $\hat{d}_{i}(t)$ and $\hat{d}_{j}(t^{\prime})$, one can define the corresponding correlation function
@@
which for a stationary state depends only on the time difference $\tau=t-t^{\prime}$. Owing to the fact that, in general, the involved operators do not commute, we may have complex-valued correlation functions even if the operators $\hat{d}_{i}$ are Hermitian [59] From Eq. (1) and the stationarity, we then further obtain that $C_{ij}^{*}(\tau)=C_{ji}(-\tau)$, where the subscript ‘*’ denotes complex conjugation. The set of correlation functions represents a correlation tensor $\underline{C}(\tau)$ so that we may introduce a matrix notation for which we accordingly have that
@@
Here, the superscript $\dagger$ denotes the Hermitian conjugate of a matrix.
Using the Wiener-Khinchin theorem, we define the power-spectrum tensor $\underline{S}(\omega)$ as
@@
$\underline{S}(\omega)=\underline{K}(\omega)+\bm{\Omega}(\omega)\cdot\mathbf{\underline{L}}~{},$ (4)
where $\underline{K}(\omega)$ is a real symmetric tensor. The quantity $\bm{\Omega}(\omega)$ is a real three-dimensional vector so that $\bm{\Omega}\cdot\mathbf{\underline{L}}\equiv\sum_{k}\Omega_{k}\cdot\underline{L}_{k}$ with $[\underline{L}_{k}]_{ij}=-\mathrm{i}\epsilon_{kij}$ being the skew-symmetric (trace-less) matrix generators of the real group of three-dimensional rotations (SO(3) group; $\epsilon_{kij}$ indicates the Levi-Civita symbol).
The first and second term on the r.h.s. of Eq. (4) correspond, respectively, to the real and imaginary part of the power-spectrum tensor, thereby uniquely identifying their associated symmetry properties. One can show the power-spectrum tensor $\underline{S}(\omega)$ is a positive semidefinite matrix (see App. A). Consequently $\underline{K}(\omega)=[\underline{S}(\omega)+\underline{S}^{\sf T}(\omega)]/2$ (the superscript ‘$\sf T$’ denotes the transpose of a matrix) is also positive semidefinite, since it is given by the sum of two positive semidefinite matrices. Given that the orthogonal transformation that diagonalizes $\underline{K}(\omega)$ is an element of SO(3), we can find a basis where in the decomposition of Eq. (4) $\underline{K}(\omega)$ is diagonal. We can write the vector $\bm{\Omega}(\omega)$ as the Fourier transform of
$\bm{\Omega}(\tau)=\frac{\mathrm{i}}{2}\langle\hat{\mathbf{d}}(t)\times\hat{\mathbf{d}}(t^{\prime})\rangle=\frac{\mathrm{i}}{2}\langle\hat{\mathbf{d}}(\tau)\times\hat{\mathbf{d}}(0)\rangle$ (5)
for which we obtain $\bm{\Omega}(-\tau)=\bm{\Omega}^{*}(\tau)$. One can interpret this as a measure for the net rotation of $\hat{\mathbf{d}}(t)$ with respect to $\hat{\mathbf{d}}(t^{\prime})$. For classical systems, a similar quantity has been associated with the occurrence of nonequilibrium dynamics [60, 61, 62], pointing to an analogous meaning for part of $\underline{S}(\omega)$ proportional to $\bm{\Omega}(\omega)$.
@@
where $\theta(t)$ denotes the Heaviside step function Once again, due to stationarity, the linear response depends only on $\tau=t-t^{\prime}$. Given that $\underline{\alpha}(\tau)$ is real, its Fourier transform fulfills what we call ‘crossing relation’, i.e. $\underline{\alpha}^{*}(\omega)=\underline{\alpha}(-\omega)$.
As long as the Fourier transform exists, the susceptibility $\underline{\alpha}(\omega)$ is analytical in the upper half of the complex frequency plane. As a consequence, it fulfills the following properties: Along the positive imaginary frequency axis, $\alpha_{ij}(\omega)$ is a real and monotonically decreasing function [63] and its real and imaginary part are, respectively, even and odd functions of $\omega$. The exchange (absorption and emission) of energy connected with our vectorial quantum-mechanical observable during a generic perturbation can be written in terms of
@@
which is proportional to the skew-Hermitian part of $\underline{\alpha}(\omega)$. Contrary to the linear susceptibility tensor, for $\omega>0$ $\underline{\alpha}_{\Im}$ is a positive definite Hermitian matrix (see App. A) and, similar to $\underline{S}$, allows for a decomposition analogous to Eq. (4). Owing to the crossing relation, we have that $\underline{\alpha}_{\Im}(-\omega)=-\underline{\alpha}^{*}_{\Im}(\omega)=-\underline{\alpha}^{\sf T}_{\Im}(\omega)$, which implies that the real part of each matrix element is odd in $\omega$, while the corresponding imaginary part is even in the same variable. As a consequence, we also have that in general $\underline{\alpha}_{\Im}(0)\neq 0$ and is skew-symmetric. The elements of $\underline{\alpha}_{\Im}(\omega)$ are directly related to the commutator appearing in Eq. (6). In particular, we obtain
@@
Since $\underline{S}$ and $\underline{\alpha}_{\Im}$ are, respectively, positive semidefinite and positive for $\omega>0$, we obtain that $\underline{S}(\omega)>\underline{S}^{*}(-\omega)=\underline{S}^{\sf T}(-\omega)$, where the inequality has to be interpreted in the sense of the Loewner order [66]. The asymmetry of Eq. (10) is shaped by the quantum nature of the system. Indeed, classically one has that $\underline{C}(\tau)\in\mathbb{R}$ implying that $\underline{S}^{*}(-\omega)=\underline{S}(\omega)$.
In the special case of a system in thermal equilibrium at the temperature $T$, i.e for the specific steady state given by
@@
[…]
Indeed, this one of of possible forms of the celebrated fluctuation-dissipation theorem (FDT) [2, 3]. Given that equilibrium is a special steady state, we can verify that Eq. (12) must explicitly fulfill Eq. (10). Specifically, here Eq.(10) can be seen as the manifestation of the asymmetry introduced by quantum fluctuations in the absorption and emission processes characterizing the system dynamics. A concrete physical example of this phenomenon is the spontaneous decay process characterizing an atom in the electromagnetic vacuum [67].
In equilibrium, the Onsager reciprocity relations [68, 69, 63] imply that the susceptibility $\underline{\alpha}$ must be symmetric so that $\underline{\alpha}_{\Im}=\underline{\alpha}_{I}$, where the subscript ‘$I$’ indicates the element-wise imaginary party of the tensor. As a consequence, the corresponding power-spectrum tensor is real and symmetric. The Onsager reciprocity relations are thus equivalent to the condition $\bm{\Omega}=0$ which indicates that in equilibrium, despite possible (quantum) fluctuations, no net rotation of $\hat{\mathbf{d}}(\tau)$ around $\hat{\mathbf{d}}(0)$ can occur [60, 61, 62]. Equivalently, from Eq. (2), we have that the correlation tensor must fulfill the crossing relation (in time) $\underline{C}^{*}(\tau)=\underline{C}(-\tau)$ (the correlation tensor becomes symmetric).
@@
[…]
Here, $\beta$ corresponds to the local equilibrium temperature of the system, often coinciding with that of its immediate surrounding. The matrix $\underline{J}(\omega)$ represents the correction to the FDT due to the system’s possible nonequilibrium dynamics. In fact, we can relate this correction to the net energy that flows between the system and the environment in the steady-state [70, 71, 72, 73, 36, 26].
The definition in Eq. (13) allows us to render the question about the validity of the local thermal equilibrium approximation more precise, i.e. as a quantitative comparison of the impact of $\underline{J}(\omega)$ on the relevant physical quantities relative to that of the FDT-like term. As long as the contribution of $\underline{J}(\omega)$ can be neglected, the local thermal equilibrium approximation provides a viable approach. For a more detailed analysis, one often has to rely on the system’s microscopic dynamics and on the specific physical quantity of interest (see also Sec. III.2). Nonetheless, we can infer a number of general properties of $\underline{J}(\omega)$. Indeed, since both $\underline{S}(\omega)$ and $\underline{\alpha}_{\Im}(\omega)$ are Hermitian, we must have that $\underline{J}(\omega)=\underline{J}^{{\dagger}}(\omega)$. Upon employing the definition of $\underline{\nu}(\omega)$, we can write

### v2

@@ Abstract
We derive and classify properties of the power-spectrum tensor for systems in general steady-states, including stationary states not necessarily corresponding to equilibrium configurations. We establish a rigorous connection between the power-spectrum tensor and other quantities that characterize these systems, providing a systematic comparison with their equilibrium counterparts. As a physical application, we investigate the problem of quantum friction, describing the contactless quantum-electrodynamic drag acting on a particle moving in close proximity to material bodies at zero temperature. Specifically, we demonstrate how including additional information about the system’s physical properties facilitates the derivation of more precise constraints on the power spectrum and its functional dependencies.
keywords:
Nonequilibrium steady states, Fluctuation-dissipation relations, Quantum friction, Fluctuational electrodynamics
articletype: Paper
1 Introduction
@@
Stochastic processes are omnipresent in engineering, physics and chemistry, so that a detailed understanding of their properties lies at the heart of many applications both, for fundamental and technological purposes. Specifically, the power spectral density is one of the most relevant quantities for studying noisy time-stationary systems: It describes how the strength (commonly the power) of different components of stochastic processes is distributed across different frequencies [1]. Within a noisy signal, this quantity allows for example to identify the dominant frequency-components, which, in turn, can be connected with the existence of specific physical properties and the existence of underlying physical phenomena. The concept of power spectral density can be generalized to multivariate or multidimensional stochastic processes where it becomes tensorial in nature and we speak of power-spectrum tensors.
@@
Among the various fundamental relations involving the power spectrum, one of the most significant is the fluctuation-dissipation theorem (FDT) [2, 3]. For systems near equilibrium, this relation links the power-spectrum tensor to the susceptibility tensor. The latter characterizes the linear response to external perturbations, thereby capturing the dynamics of the system in the vicinity of its unperturbed state. Specifically, in the FDT both the power-spectrum and susceptibility tensors are defined as expectation values evaluated over the system’s thermal equilibrium state. The theorem is related to the detailed balance between the different spontaneous energy-exchange processes that take place in the system [4, 5]. Probably one of the most remarkable feature of the FDT is its generality: As long as it is possible to define the system’s power spectrum and susceptibility for the thermal equilibrium state, the details of the underlying microscopic dynamics do not matter. Whether the dynamics is linear or nonlinear, whether it is quantum or classical, whether it is Markovian or not [4, 1, 6], in the FDT the power spectrum is always proportional to the susceptibility via a known factor that depends on temperature. Therefore, it is not at all surprising that the theorem is applied in numerous and very diverse physical systems [7, 8, 9, 10, 11, 12, 13], sometimes even transcending its range of applicability. In fact, despite the fact that the FDT is valid close to equilibrium, i.e. when the power spectrum and the susceptibility are evaluated at equilibrium, it has been considered also for systems with states being generically out of thermal equilibrium [14, 15, 16, 17, 18, 19, 20] for which, technically speaking, it might become inaccurate. Arguments for the validity of this approach have typically been based on the so-called local thermal equilibrium (LTE) approximation. Within the LTE approximation, it is assumed that the correlation length and correlation time associated with the fluctuating dynamics are so short that the system’s spatially separated components are (locally) in thermal equilibrium with their immediate surroundings [21, 22, 23]. In a corse graining approach this usually allows for the definition of local intensive variables like the temperature which, contrary to a global equilibrium state, might now depend on the time and on the position within the system. The quality of this approximation depends strongly on the specific system, its physical composition and parameters [24, 25] as well as on the involved length scales [21, 23], thus rendering it at least questionable in a number of cases.
@@
In this work, we examine the power-spectrum tensor and its fundamental properties, generalizing its connection to other physical quantities by evaluating expectation values over a nonequilibrium steady state (NESS) rather than an equilibrium state. In such stationary states, a system’s stochastic properties are translation invariant with respect to time, despite the fact that the system can be far from equilibrium. In recent years, such stationary states have attracted particular attention [26, 27, 5, 28, 29, 30, 31, 32, 33] largely motivated by the research on fluctuation theorems more general than FDT and thermodynamic uncertainty relations [6, 34, 33, 35, 36, 37, 38, 39, 40, 41, 42]. These results establish connections between different physical quantities and lead to bounding inequalities with a broad range of validity, even beyond equilibrium. To exist, the NESS must thus be sustained by one or several external agents such as externally maintained temperature gradients, chemical potential differences, or even mechanical drives that act on the system. In general, this can break the aforementioned detailed balance and, therefore, lead to violations of the FDT [5].
@@
While we address nonequilibrium dynamics, our approach also includes the equilibrium case as a special limit. Specifically, we limit the assumptions about the microscopic dynamics of the system, thereby keeping the level of generality similar to that of the FDT. The stationarity of the NESS still allows us to draw a number of conclusions about the form and the mathematical properties of the power spectrum, allowing to obtain specific expressions that can be related to the physical characteristics of the underlying stochastic dynamics. These results can be useful to determine and analyze physical quantities relying on the power spectrum and to verify the consistency of expressions obtained using more detailed descriptions.
@@
As an example for the validity of our results, we further discuss the properties of the power-spectrum tensor in the case of the nonconservative viscous motion of a particle in the quantum electromagnetic vacuum [15, 43, 44, 45, 46]. Specifically, we address quantum friction, i.e. the drag force acting on a neutral particle driven with constant velocity in the empty space close to a generic arrangement of bodies. Still keeping the microscopic dynamics unspecified, we show how additional information about the physics of the system strongly constrains the functional behavior and the structure of the power-spectrum tensor.
@@
The paper is organized as follows. In Sec. 2, we consider the power-spectral density in the steady-state of a multidimensional stochastic process. We analyze its mathematical properties and how they relate to the physics of the system, establishing a connection to other quantities characterizing the system like its susceptibility. In Sec. 3, we specify our analysis to the case of quantum friction and show how the thus increased physical information about the system allows us to derive additional relations. To confirm our results and illustrate the advantages as well as the simplifications introduced by our general approach, we consider here an exactly solvable model for describing the internal dynamics of the particle. Finally, in Sec. 4 we summarize and discuss our findings.
2 The power-spectrum tensor in steady-state systems
@@
We consider a generic, potentially driven quantum system. Its dynamics are generated by a Hamiltonian $\hat{H}_{t}$ that may explicitly depend on time. Our focus here is on a quantum system with a large number of degrees of freedom. It is thought, in general, as being composed of a collection of interacting subsystems: On its own, each of the quantum subsystems is open, i.e. able to exchange energy and other physical quantities with the remaining parts of the system, representing its environment [47, 48] (see Fig. 1a). Therefore, the reduced dynamics of a subsystem is in general non-Hermitian [47, 48, 49]. Contrary to typical approaches, however, we will refrain from deriving any master equation, avoiding the Markov and other approximations which, depending on the specific circumstances, can be problematic for the description of quantum systems [4, 50, 51]. Our framework remains independent of microscopic details and assumes only that the system has reached a unique steady state $\hat{\rho}$, characterized by the condition $\hat{\rho},\hat{H}_{t}]=0$. Again, since energy can enter the system due to the drive, we would like to note explicitly that the time independent density operator $\hat{\rho}$ does not necessarily correspond to an equilibrium state and it can describe a system far from equilibrium.
Refer to captionab
Figure 1: a: Sketch of the generic configuration analyzed in this work. A large, potentially driven quantum system composed of a subsystem interacting with an environment. The environment itself can be resulting from the interaction of different subsystems. b: A possible physical implementation of this generic scheme: An externally driven atom coupled to a complex quantum environment, emerging from the self-consistent dynamics of the electromagnetic field in the presence of material bodies. The energy entering the system through the drive can be dissipated away in form of electromagnetic radiation and heat within the material, allowing the systems to reach a nonequilibrium steady state.
@@
Let $\hat{d}_{i}$ be the components of a Hermitian vector-operator $\hat{\mathbf{d}}$, describing physical quantities belonging to one of the interacting subsystems. The time-evolution in the Heisenberg picture of each component, $\hat{d}_{i}(t)$, is induced according to $\hat{H}_{t}$ and we assume that all the components exhibit the same time-reversal symmetry. Due to the interaction between the subsystems, the time-evolved operator $\hat{d}_{i}(t)$ encodes the global dynamics, reflecting the potentially high degree of entanglement between the system’s various degrees of freedom. For our purposes, we consider a three-dimensional vector but generalizations to different dimensions are possible. The vector-operator $\hat{\mathbf{d}}$ either corresponds to a three-dimensional observable, or a quantity that characterizes a multivariate stochastic process as for example in three-particle systems. For simplicity, in the subsequent discussion, we will refer to the former case but most of the results can be transferred to the latter case upon adequate replacements of the relevant quantities. As a specific example, $\hat{\mathbf{d}}$ can describe a particle’s electric dipole moment ($i=x,y,z$), interacting with a complex electromagnetic environment including objects comprised of dispersive and dissipative materials (see Fig. 1b and Sec. 3).
@@
Since the system is stationary, the expectation value of $\hat{\mathbf{d}}(t)$ over $\hat{\rho}$ is constant. Without loss of generality, we assume it to be zero  [^1] This condition can always be realized if we consider instead of $\hat{d}_{i}(t)$ the operator $\hat{\tilde{d}}_{i}(t)=\hat{d}_{i}(t)-\langle\hat{d}_{i}\rangle$., i.e. $\langle\hat{d}_{i}(t)\rangle\equiv\mathrm{tr}[\hat{d}_{i}(t)\hat{\rho}]=\mathrm{tr}[\hat{d}_{i}\hat{\rho}]=0$, where $\mathrm{tr}$’ indicates the trace over a generic set of basis states of the whole system’s Hilbert space. Given $\hat{d}_{i}(t)$ and $\hat{d}_{j}(t^{\prime})$, one can define the corresponding correlation function
@@
which, for a stationary state, depends only on the time difference $\tau=t-t^{\prime}$. Owing to the fact that, in general, the involved operators do not commute, we may have complex-valued correlation functions even if the operators $\hat{d}_{i}$ are Hermitian  [^2] In principle, we can avoid this by symmetrizing the product in the definition of $C_{ij}(\tau)$ (see for example [52, 53]).. From Eq. (1) and the stationarity, we then further obtain that $C_{ij}^{*}(\tau)=C_{ji}(-\tau)$, where the subscript ‘*’ denotes complex conjugation. The set of correlation functions represents a correlation tensor $\underline{C}(\tau)$ so that we may introduce a matrix notation for which we accordingly have that
@@
Here, the superscript $\dagger$’ denotes the Hermitian conjugate.
Relying on the Wiener-Khinchin theorem [1], the power-spectrum tensor $\underline{S}(\omega)$ is given by
@@
$\underline{S}(\omega)=\underline{K}(\omega)+\boldsymbol{\Omega}(\omega)\cdot\mathbf{\underline{L}}~,$ (4)
where $\underline{K}(\omega)$ is a real symmetric matrix. The quantity $\boldsymbol{\Omega}(\omega)$ is a real three-dimensional vector so that $\boldsymbol{\Omega}\cdot\mathbf{\underline{L}}\equiv\sum_{k}\Omega_{k}\cdot\underline{L}_{k}$. Here $\underline{L}_{k}]_{ij}=-\mathrm{i}\epsilon_{kij}$ are the skew-symmetric (trace-less) matrix generators of the real group of three-dimensional rotations (SO(3) group; $\epsilon_{kij}$ indicates the Levi-Civita symbol).
The first and second term on the r.h.s. of Eq. (4) correspond, respectively, to the real and imaginary part of the power-spectrum tensor, thereby uniquely identifying their associated symmetry properties. The power-spectrum tensor $\underline{S}(\omega)$ is a positive semidefinite matrix (see App. A). Consequently $\underline{K}(\omega)=[\underline{S}(\omega)+\underline{S}^{\sf T}(\omega)]/2$ (the superscript $\sf T$’ denotes the transpose operation) is also positive semidefinite, since it is given by the sum of two positive semidefinite matrices. Given that the orthogonal transformation that diagonalizes $\underline{K}(\omega)$ is an element of SO(3), we can find a basis where, in the decomposition of Eq. (4), $\underline{K}(\omega)$ is diagonal. We can write the vector $\boldsymbol{\Omega}(\omega)$ as the Fourier transform of
$\boldsymbol{\Omega}(\tau)=\frac{\mathrm{i}}{2}\langle\hat{\mathbf{d}}(t)\times\hat{\mathbf{d}}(t^{\prime})\rangle=\frac{\mathrm{i}}{2}\langle\hat{\mathbf{d}}(\tau)\times\hat{\mathbf{d}}(0)\rangle$ (5)
for which we obtain $\boldsymbol{\Omega}(-\tau)=\boldsymbol{\Omega}^{*}(\tau)$. One can then interpret $\boldsymbol{\Omega}$ as a measure for the net rotation of $\hat{\mathbf{d}}(t)$ with respect to $\hat{\mathbf{d}}(t^{\prime})$. For classical systems, a similar quantity has been associated with the occurrence of nonequilibrium dynamics [54, 55, 56], pointing to an analogous meaning for the part of $\underline{S}(\omega)$ that is proportional to $\boldsymbol{\Omega}(\omega)$.
@@
where $\theta(t)$ denotes the Heaviside step function. Once again, due to stationarity, the linear response depends only on $\tau=t-t^{\prime}$. Given that $\underline{\alpha}(\tau)$ is real, its Fourier transform fulfills what we call ‘crossing relation’, i.e. $\underline{\alpha}^{*}(\omega)=\underline{\alpha}(-\omega)$.
As long as the Fourier transform exists, the Paley–Wiener theorem [57] implies that the susceptibility $\underline{\alpha}(\omega)$ is analytical in the upper half of the complex frequency plane. As a consequence, it fulfills the following properties: Along the positive imaginary frequency axis, $\alpha_{ij}(\omega)$ is a real and monotonically decreasing function [53] and its real and imaginary part are, respectively, even and odd functions of $\omega$. The exchange (absorption and emission) of energy connected with our vectorial quantum-mechanical observable due to a generic perturbation can be written in terms of
@@
which is proportional to the skew-Hermitian part of $\underline{\alpha}(\omega)$. Contrary to the linear susceptibility tensor, $\underline{\alpha}_{\Im}$ is a positive definite Hermitian matrix for $\omega>0$ (see App. A) and allows for a decomposition analogous to Eq. (4). Owing to the crossing relation, we have that $\underline{\alpha}_{\Im}(-\omega)=-\underline{\alpha}^{*}_{\Im}(\omega)=-\underline{\alpha}^{\sf T}_{\Im}(\omega)$, which implies that the real part of each matrix element is odd in $\omega$, while the corresponding imaginary part is even in the same variable. As a consequence, we also have that in general $\underline{\alpha}_{\Im}(0)\neq 0$ and is skew-symmetric. In connection with the previous properties, we can show that in general the symmetric part of despite $\underline{\alpha}(\mathrm{i}\xi)$ must be positive definite. Indeed we have (see App. A)
$\frac{\underline{\alpha}(\mathrm{i}\xi)+\underline{\alpha}^{\sf T}(\mathrm{i}\xi)}{2}=\frac{2}{\pi}\mathrm{Re}\int_{0}^{\infty}\mathrm{d}\omega\frac{\omega}{\omega^{2}+\xi^{2}}\underline{\alpha}_{\Im}(\omega)~.$ (8)
The elements of $\underline{\alpha}_{\Im}(\omega)$ are directly related to the commutator appearing in Eq. (6). In particular, we obtain
@@
Since $\underline{S}$ and $\underline{\alpha}_{\Im}$ are, respectively, positive semidefinite and positive for $\omega>0$, we obtain that $\underline{S}(\omega)>\underline{S}^{*}(-\omega)=\underline{S}^{\sf T}(-\omega)$, where the inequality has to be interpreted in the sense of the Loewner order [59]. The asymmetry of Eq. (11) is shaped by the quantum nature of the system. Indeed, classically one has that $\underline{C}(\tau)\in\mathds{R}$ implying that $\underline{S}^{*}(-\omega)=\underline{S}(\omega)$.
In the special case of the steady state being thermal equilibrium at the temperature $T$, i.e when the system is described by a time-independent Hamiltonian $\hat{H}_{0}$ and
@@
$\underline{S}(\omega)=\frac{\hbar}{\pi}\frac{\underline{\alpha}_{\Im}(\omega)}{1-e^{-\beta\hbar\omega}}\quad\text{(Equilibrium state)}.$ (13)
Indeed, this represents one of the possible forms of the FDT. Given that equilibrium is a special steady state, we can verify that Eq. (13) must explicitly fulfill Eq. (11). Specifically, here Eq.(11) can be seen as the manifestation of the asymmetry introduced by quantum fluctuations in the absorption and emission processes characterizing the system dynamics. A concrete physical example of this phenomenon is the spontaneous decay process characterizing an atom in the electromagnetic vacuum [60].
Whenever the Onsager reciprocity relations [61, 62, 63, 53] are valid, they imply that, when averaged over the equilibrium state, the susceptibility $\underline{\alpha}$ must be symmetric so that $\underline{\alpha}_{\Im}\equiv\underline{\alpha}_{I}$ is real, where the subscript ‘$I$’ indicates the element-wise imaginary party of the tensor. Considering Eq. (13), one hat that the corresponding power-spectrum tensor is also real and, taking into account Eq. (4), we can deduce that it is symmetric. For reciprocal systems close to equilibrium, the Onsager reciprocity relations are thus equivalent to the condition $\boldsymbol{\Omega}=0$ which indicates that for an equilibrium state, despite possible (quantum) fluctuations, no net rotation of $\hat{\mathbf{d}}(\tau)$ around $\hat{\mathbf{d}}(0)$ can occur [54, 55, 56]. Equivalently, from Eq. (2), we have that the correlation tensor must fulfill the crossing relation (in time) $\underline{C}^{*}(\tau)=\underline{C}(-\tau)$ (the correlation tensor becomes symmetric).
@@
$\underline{S}(\omega)=\frac{\hbar}{\pi}\frac{\underline{\alpha}_{\Im}(\omega)}{1-e^{-\beta\hbar\omega}}+\underline{J}(\omega)\quad\text{(NESS)}~.$ (14)
Here, $\beta$ corresponds to the local equilibrium temperature of the system, often coinciding with that of its immediate surrounding. The matrix $\underline{J}(\omega)$ represents the correction to the FDT due to the system’s possible far-from-equilibrium dynamics, going beyond the typical framework provided by linear-response theory. In fact, we can relate this correction to the net energy that flows between the system and the environment in the NESS [64, 65, 66, 67, 40, 23].
The definition in Eq. (14) allows us to render the question about the validity of the local thermal equilibrium approximation more precise, i.e. as a quantitative comparison of the impact of $\underline{J}(\omega)$ on the relevant physical quantities relative to that of the FDT-like term. As long as the contribution of $\underline{J}(\omega)$ can be neglected, the local thermal equilibrium approximation provides a viable approach. For a more detailed analysis, one often has to rely on the system’s microscopic dynamics and on the specific physical quantity of interest (see also Sec. 3.2). Nonetheless, we can infer a number of general properties of $\underline{J}(\omega)$. Indeed, since both $\underline{S}(\omega)$ and $\underline{\alpha}_{\Im}(\omega)$ are Hermitian, we must have that $\underline{J}(\omega)=\underline{J}^{{\dagger}}(\omega)$. Upon employing the definition of $\underline{\nu}(\omega)$, we can write